OpenStudy (cheesecakekitten):

Solve: 5x^2 - 45 = 0 A) 9 B) 3√5 C) ±3√5 D) ±3

OpenStudy (cheesecakekitten):

I can see why broonie cat was angru with opestud

OpenStudy (isaiah.feynman):

Sorry.

OpenStudy (cheesecakekitten):

halp plz

OpenStudy (cheesecakekitten):

IDK HOW ;-;

OpenStudy (isaiah.feynman):

Solve it algebraically.

OpenStudy (isaiah.feynman):

\[5x^{2}-45 = 0\]

OpenStudy (isaiah.feynman):

Add 45 to both sides, divide both sides by 5, and square both sides to get \[x = \pm \sqrt{\frac{45 }{ 5 }} = \pm \sqrt{9} = \pm3\]

OpenStudy (cheesecakekitten):

Okay...

OpenStudy (qwertty123):

Cheesecake... did you get the other question answered?

OpenStudy (qwertty123):

BTW @Isaiah.Feynman please don't do all the work for the user.. That person should be helping you too.. :)

OpenStudy (isaiah.feynman):

Sure thing.

alones (alones):

@Isaiah.Feynman Ddi the right thing.. hope kitted attention ;3

OpenStudy (cheesecakekitten):

@AloneS @Isaiah.Feynman

OpenStudy (cheesecakekitten):

hallllp

alones (alones):

@CheesecakeKitten make a post i will try to help i'll try